Dave averaged 30 miles per hour on his bike trip to a friend's house. His return trip took an hour longer, as he averaged 20 miles per hour. What was his one-way distance to his friends?
d = s * t

Answer by msaxby(1)

Let t = time to the friends house.
Since d = s *t, 30*t = distance to the friend's house.
and 20*(t+1) = the distance back from the friend's house.
SO: 30*t = 20*(t+1)
30*t = 20*t + 20
10*t = 20
t = 2 (hours).
Since Dave averaged 30 miles per hour, it is 30*2 miles to his friend's house or 60 miles.
(Check: Averaging 20 mph would take 60 / 20 or 3 hours - that's 1 hour longer.)